OpenStudy (anonymous):

What is the equation in standard form of the line which passes through (6, 3) and has a slope of ? 2x + 3y = 3 2x – 3y = 3 3x + 2y = 3 3x – 2y = 3

OpenStudy (anonymous):

do you know what the standard form of a linear equation looks like?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

y=mx+b m is the slope and b is the y intercept. so.. is your question incomplete or are u supposed to use the given equations to compute the slope?

OpenStudy (anonymous):

the slope is 2/3

jimthompson5910 (jim_thompson5910):

Hint: In Ax + By = C the slope is -A/B In Ax - By = C the slope is A/B

OpenStudy (anonymous):

if you know the slope and the point it passes through you can use the point slope formula to figure out the equation of the line. y-y1 = m(x-x1)

OpenStudy (anonymous):

so y - 3 = 2/3(x - 6)

OpenStudy (anonymous):

can you solve from there and rearrange to the right form?

OpenStudy (anonymous):

I can't get the fraction part figured out

jimthompson5910 (jim_thompson5910):

Hint: start with y - 3 = 2/3(x - 6) and multiply everything (but the stuff in the parenthesis) by 3 to get 3y - 9 = 2(x-6)

OpenStudy (anonymous):

I got 2x-3y=3.

jimthompson5910 (jim_thompson5910):

you are correct

OpenStudy (anonymous):

Awesome. Thanks muchos

jimthompson5910 (jim_thompson5910):

yw